# AuroraCap: Efficient, Performant Video Detailed Captioning and a New Benchmark

### v1

@@ Appendix D Ablation Studies
Figure D10: Comparison between performance with and without the inclusion of full first-frame visual tokens during inference on VDC. As the visual token kept ratio varies, the blue curve indicates performance with slowfast inference, while the red curve represents performance without slowfast inference.
Appendix E Detailed Training Settings
We use CLIP ViT-H  [^5] Huggingface Model: https://huggingface.co/apple/DFN5B-CLIP-ViT-H-14-378, Vicuna-1.5-7B  [^6] Huggingface Model: https://huggingface.co/lmsys/vicuna-7b-v1.5-16k as the initialization of AuroraCap-7B. Training hyper-parameters for both stages are shown in Table E6. For visual data preprocessing, we resize each image or video frame to the short side of $378$ while keeping original aspect. Instead of doing center crop, we conduct bilinear position embedding interpolation for each training sample. For video data, we extract frames at $2$ FPS uniformly. For token merging, we use constant schedule for each blocks.
Table E3: Summary of datasets used for training AuroraCap in Pretraining stage.
Task # Sample Dataset
Image Captioning 1.3M LAION-CC-SBU-595K Liu et al. (2024b), ShareGPT4V Chen et al. (2023a), ALLaVA-Caption-LAION-4V Chen et al. (2024a), ALLaVA-Caption-VFLAN-4V Chen et al. (2024a), DenseFusion Li et al. (2024d)
Table E4: Summary of datasets used for training AuroraCap in Vision stage. For classification, Reasoning, VQA, and Generation tasks, we adopt the dataset processed by M3IT Li et al. (2023c) to fit the training objective of language models.
Task # Sample Dataset
Captioning 1,925K ShareGPT4V-PT Chen et al. (2023a), TextCaps Sidorov et al. (2020), Image-Paragraph-Captioning Krause et al. (2017)
Object-centric 438K COST Jain et al. (2023), ChatterBox Tian et al. (2024), V* Wu & Xie (2023)
Classification 238K COCO-GOI Lin et al. (2014), COCO-Text Veit et al. (2016), ImageNet Russakovsky et al. (2015), COCO-ITM Lin et al. (2014),
e-SNLI-VE Kayser et al. (2021), Mocheg Yao et al. (2023), IQA Duanmu et al. (2021)
Reasoning 100K CLEVR Johnson et al. (2017), NLVR Suhr et al. (2017), VCR Zellers et al. (2019), VisualMRC Tanaka et al. (2021), Winoground Thrush et al. (2022)
VQA 3,518K VQA v2 Goyal et al. (2017), Shapes VQA Andreas et al. (2016), DocVQA Mathew et al. (2021), OK-VQA Marino et al. (2019),
Text-VQA Singh et al. (2019), OCR-VQA Mishra et al. (2019), A-OK-VQA Schwenk et al. (2022), ScienceQA Lu et al. (2022)
ST-VQA Biten et al. (2019), ViQuAE Lerner et al. (2022), LLaVA-OneVision Li et al. (2024a)
Generation 145K Visual Storytelling Huang et al. (2016), Visual Dialog Das et al. (2017), Multi30k Elliott et al. (2016)
Chinese 193K COCO-Caption CN Li et al. (2019), Flickr-8k-Caption CN Li et al. (2016), multimodal Chat Zheng et al. (2021), FM-IQA Gao et al. (2015), ChineseFoodNet Chen et al. (2017)
Table E5: Summary of datasets used for training AuroraCap in Language stage.
Task # Sample Dataset
Image Captioning 1,779K ShareGPT4V Chen et al. (2023a), ALLaVA-Caption-LAION-4V Chen et al. (2024a), ALLaVA-Caption-VFLAN-4V Chen et al. (2024a), DenseFusion Li et al. (2024d), FaceCaption Dai et al. (2024a)
Video Captionin 1,659K MiraData Ju et al. (2024), LLaVA-Hound Zhang et al. (2024c), ShareGPT4Video Chen et al. (2024d), Private Data
Image Instruction 9,742K LLaVA-Mix-665K Liu et al. (2023), LVIS-Instruct4V Wang et al. (2023), ALLaVA-Instruct-LAION-4V Chen et al. (2024a), ALLaVA-Instruct-VFLAN-4V Chen et al. (2024a), Cambrian Tong et al. (2024a), M4-Instruct Liu et al. (2024a)
Video Instruction 268k LLaVA-Hound Zhang et al. (2024c), ShareGPT4Video Chen et al. (2024d)
Language-only 143K Evol-Intruct-GPT4-Turbo-143K Chen et al. (2024a)
Table E6: Training hyper-parameters for AuroraCap.
@@ Appendix E Detailed Training Settings
optimizer AdamW Loshchilov & Hutter (2017)
@@
Appendix F Evaluation Benchmarks and Settings
We list all the hyper-parameters and prompt used for evaluation as shown in Table F7. For our proposed VDC, we show the settings as following with the max number of the tokens of 1,024:
@@ Appendix F Evaluation Benchmarks and Settings
Table F7: Evaluation settings summary for each benchmarks. For all benchmarks we set temperature, top p, number of beams to 0, 0, 1 respectively.
@@
Flickr Plummer et al. (2015) 31,784 64 Provide a one-sentence caption for the provided image.
NoCaps Agrawal et al. (2019) 4,500 64 Provide a one-sentence caption for the provided image.
COCO-Cap Lin et al. (2014) 5,000 64 Provide a one-sentence caption for the provided image.
ChartQA Masry et al. (2022) 2,500 16 Answer the question with a single word.
DocVQA Mathew et al. (2021) 5,349 32 Answer the question using a single word or phrase.
TextCaps Sidorov et al. (2020) 3,166 64 Provide a one-sentence caption for the provided image.
GQA Hudson & Manning (2019) 12,578 16 Answer the question using a single word or phrase.
POPE Li et al. (2023h) 9,000 128 Answer the question using a single word or phrase.
MMMU Yue et al. (2023) 900 16 Answer with the option letter from the given choices directly.
@@
VQAv2 Goyal et al. (2017) 214,354 16 Answer the question using a single word or phrase.
MSR-VTT Xu et al. (2016) 1,000 64 Provide a one-sentence caption for the provided video.
VATEX Wang et al. (2019) 4,478 64 Provide a brief single-sentence caption for the last video below.
MSVD-QA Xu et al. (2017) 1,161 64 Answer the question using a single word or phrase.
ActivityNet-QA Yu et al. (2019) 8,000 64 Answer the question using a single word or phrase.
MSRVTT-QA Xu et al. (2017) 6,513 64 Answer the question using a single word or phrase.
iVQA Yang et al. (2021) 6,000 64 Answer the question using a single word or phrase.
Appendix G Limitations
Table G8: Comparison AuroraCap with LLM-based SoTA methods on visual question answering benchmarks under zero-shot setting.
@@ Appendix G Limitations
Table G9: Limitation in terms of OCR capability compared with LLaVA models. Appendix F shows the introduction and metrics of each benchmark.
@@
We evaluate AuroraCap on various visual question answering benchmarks as shown in Table G8. Since the performance of the VQA task heavily depends on the performance of the LLM, we chose the same LLM for a fair comparison. Also, due the the limitation of OCR-related samples in training dataset, AuroraCap does not perform well in OCR at current stage as shown in Table G9.
Appendix H VDC Generation Prompt Template
Following Ju et al. (2024), we utilize LLM to generate structured detailed captions. Given an input video, LLM return five detailed captions, including camera caption, short caption, background caption, main object caption and detailed caption for the entire video guided by our designed prompt template. The complete prompt is shown as followings:
@@ Appendix H VDC Generation Prompt Template
Appendix I Question-answer Pairs Generation Prompt Template of VDCscore
@@
Appendix J More Statistics Information of VDC
Figure J11 indicates the word distribution of the structured captions in VDC. Table J10 illustrates the distribution of the question categories we provided for VDCscore calculation.
Figure J11: Word cloud of different structured captions in VDC, showing the diversity.
Table J10: Distribution of the question categories in VDCscore.
@@
Appendix K More Statistics Information of VDCscore
We generate a total of 96,902 question-answer pairs for VDC, with an average of 18.87 pairs per detailed caption. As depicted in Figure K13, each section of the structured captions includes a similar number of question-answer pairs. Additionally, Figure K13 presents the distribution of question types generated for VDC. To enhance the evaluation of detailed captions, we configure all questions as open-ended. Environment & Scene encompasses inquiries about location, environment, atmosphere, scene, and time. Character & Object focuses on entities within the caption, such as people, animals, plants, and objects, while Attribute & Relation examines their attributes and interrelations. Intent & Outcomes addresses deeper interpretative questions regarding methods, purposes, reasons, and outcomes. We further analyze the distribution of generated question types within structured captions. Main object captions predominantly feature Character & Object questions, whereas Environment & Scene questions are more prominent in background captions, and Camera questions constitute a larger proportion in camera.
Figure K12: Question-answer pairs proportaion in structured captions.
Figure K13: Distribution of question type in VDCscore.
Appendix L Calculation of Elo Ranking
In this section, we present the methodology for evaluating and ranking  AuroraCapand various models using the Elo rating system as shown in Figure L14. The parameters used in the simulation are summarized in Table L11.
Table L11: Elo parameter setting.
@@
Figure L14: GUI screen for Elo ranking.
Appendix M Case Study
@@
Figure M15: Video example of MSR-VTT Xu et al. (2016) benchmark.
Figure M16: Video example of VATEX Wang et al. (2019) benchmark.
Figure M17: Video example.
Figure M18: Video example.
Figure M19: Video example.
Figure M20: Video example.
Figure M21: Video example.
Figure M22: Video example.

### v2

@@ 10 Ablation Studies
11 Detailed Training Settings
We use CLIP ViT-H ***Huggingface Model: https://huggingface.co/apple/DFN5B-CLIP-ViT-H-14-378, Vicuna-1.5-7B ***Huggingface Model: https://huggingface.co/lmsys/vicuna-7b-v1.5-16k as the initialization of AuroraCap-7B. Training hyper-parameters for both stages are shown in Table 12. For visual data preprocessing, we resize each image or video frame to the short side of $378$ while keeping original aspect. Instead of doing center crop, we conduct bilinear position embedding interpolation for each training sample. For video data, we extract frames at $2$ FPS uniformly. For token merging, we use constant schedule for each blocks.
Table 12: Training hyper-parameters for AuroraCap.
@@ 11 Detailed Training Settings
optimizer AdamW [83]
@@
Table 13: Summary of datasets used for training AuroraCap in Pretraining stage.
Task # Sample Dataset
Image Captioning 1.3M LAION-CC-SBU-595K [82], ShareGPT4V [18], ALLaVA-Caption-LAION-4V [15], ALLaVA-Caption-VFLAN-4V [15], DenseFusion [68]
Table 14: Summary of datasets used for training AuroraCap in Vision stage. For classification, Reasoning, VQA, and Generation tasks, we adopt the dataset processed by M3IT [65] to fit the training objective of language models.
Task # Sample Dataset
Captioning 1,925K ShareGPT4V-PT [18], TextCaps [111], Image-Paragraph-Captioning [55]
Object-centric 438K COST [48], ChatterBox [123], V* [135]
Classification 238K COCO-GOI [78], COCO-Text [128], ImageNet [108], COCO-ITM [78],
e-SNLI-VE [53], Mocheg [149], IQA [33]
Reasoning 100K CLEVR [50], NLVR [116], VCR [154], VisualMRC [118], Winoground [122]
VQA 3,518K VQA v2 [39], Shapes VQA [4], DocVQA [92], OK-VQA [90],
Text-VQA [113], OCR-VQA [94], A-OK-VQA [109], ScienceQA [84]
ST-VQA [9], ViQuAE [59], LLaVA-OneVision [60]
Generation 145K Visual Storytelling [44], Visual Dialog [28], Multi30k [35]
Chinese 193K COCO-Caption CN [70], Flickr-8k-Caption CN [69], multimodal Chat [167], FM-IQA [37], ChineseFoodNet [22]
Table 15: Summary of datasets used for training AuroraCap in Language stage.
Task # Sample Dataset
Image Captioning 1,779K ShareGPT4V [18], ALLaVA-Caption-LAION-4V [15], ALLaVA-Caption-VFLAN-4V [15], DenseFusion [68], FaceCaption [26]
Video Captionin 1,659K MiraData [51], LLaVA-Hound [159], ShareGPT4Video [19], Private Data
Image Instruction 9,742K LLaVA-Mix-665K [80], LVIS-Instruct4V [132], ALLaVA-Instruct-LAION-4V [15], ALLaVA-Instruct-VFLAN-4V [15], Cambrian [124], M4-Instruct [81]
Video Instruction 268k LLaVA-Hound [159], ShareGPT4Video [19]
Language-only 143K Evol-Intruct-GPT4-Turbo-143K [15]
12 Evaluation Benchmarks and Settings
We list all the hyper-parameters and prompt used for evaluation as shown in Table 16. For our proposed VDC, we show the settings as following with the max number of the tokens of 1,024:
@@ 12 Evaluation Benchmarks and Settings
Table 16: Evaluation settings summary for each benchmarks. For all benchmarks we set temperature, top p, number of beams to 0, 0, 1 respectively.
@@
Flickr [102] 31,784 64 Provide a one-sentence caption for the provided image.
NoCaps [1] 4,500 64 Provide a one-sentence caption for the provided image.
COCO-Cap [78] 5,000 64 Provide a one-sentence caption for the provided image.
ChartQA [91] 2,500 16 Answer the question with a single word.
DocVQA [92] 5,349 32 Answer the question using a single word or phrase.
TextCaps [111] 3,166 64 Provide a one-sentence caption for the provided image.
GQA [45] 12,578 16 Answer the question using a single word or phrase.
POPE [74] 9,000 128 Answer the question using a single word or phrase.
MMMU [153] 900 16 Answer with the option letter from the given choices directly.
@@
VQAv2 [39] 214,354 16 Answer the question using a single word or phrase.
MSR-VTT [140] 1,000 64 Provide a one-sentence caption for the provided video.
VATEX [133] 4,478 64 Provide a brief single-sentence caption for the last video below.
MSVD-QA [138] 1,161 64 Answer the question using a single word or phrase.
ActivityNet-QA [152] 8,000 64 Answer the question using a single word or phrase.
MSRVTT-QA [138] 6,513 64 Answer the question using a single word or phrase.
iVQA [145] 6,000 64 Answer the question using a single word or phrase.
13 Additional Evaluation Metrics Analysis
Refer to caption Refer to caption Refer to caption
Figure 21: Correlation between the CLIP evaluation scores, different number of QA pairs used in VDCscore and human elo ranking on VDC.
In this section, we make the analysis about more evaluation metrics on VDC other than VDCscore. As shown in Figure 21, we found that the correlation between CLIP score and human ratings is low (0.39), which may be due to the fact that, in a lengthy caption, there may be only one or two fine-grained errors that the CLIP text encoder lacks the capability to detect. Meanwhile, we test the impact of different numbers of QA pairs on the results when calculating VDCscore. As shown in the Figure 21, more QA pairs produced more robust test results. VDCscore shows over 0.86 of pearson correlation with human when using 20 QA pairs.
14 Limitations
We evaluate AuroraCap on various visual question answering benchmarks as shown in Table 17. Since the performance of the VQA task heavily depends on the performance of the LLM, we chose the same LLM for a fair comparison. Also, due the the limitation of OCR-related samples in training dataset, AuroraCap does not perform well in OCR at current stage as shown in Table 18.
Table 17: Comparison AuroraCap with LLM-based SoTA methods on visual question answering benchmarks under zero-shot setting.
@@ 14 Limitations
Table 18: Limitation in terms of OCR capability compared with LLaVA models. Appendix 12 shows the introduction and metrics of each benchmark.
@@
15 VDC Generation Prompt Template
Following [51], we utilize LLM to generate structured detailed captions. Given an input video, LLM return five detailed captions, including camera caption, short caption, background caption, main object caption and detailed caption for the entire video guided by our designed prompt template. The complete prompt is shown as followings:
@@ 15 VDC Generation Prompt Template
16 Question-answer Pairs Generation Prompt Template of VDCscore
@@
17 More Statistics Information of VDC
Figure 22 indicates the word distribution of the structured captions in VDC. Table 19 illustrates the distribution of the question categories we provided for VDCscore calculation.
[Camera Caption]Refer to caption [Short Caption]Refer to caption [Background Caption]Refer to caption
[Main Object Caption]Refer to caption [Detailed Caption]Refer to caption [Total]Refer to caption
Figure 22: Word cloud of different structured captions in VDC, showing the diversity.
Table 19: Distribution of the question categories in VDCscore.
@@ 17 More Statistics Information of VDC
18 More Statistics Information of VDCscore
We generate a total of 96,902 question-answer pairs for VDC, with an average of 18.87 pairs per detailed caption. As depicted in Figure 24, each section of the structured captions includes a similar number of question-answer pairs. Additionally, Figure 24 presents the distribution of question types generated for VDC. To enhance the evaluation of detailed captions, we configure all questions as open-ended. Environment & Scene encompasses inquiries about location, environment, atmosphere, scene, and time. Character & Object focuses on entities within the caption, such as people, animals, plants, and objects, while Attribute & Relation examines their attributes and interrelations. Intent & Outcomes addresses deeper interpretative questions regarding methods, purposes, reasons, and outcomes. We further analyze the distribution of generated question types within structured captions. Main object captions predominantly feature Character & Object questions, whereas Environment & Scene questions are more prominent in background captions, and Camera questions constitute a larger proportion in camera.
Figure 23: Question-answer pairs proportaion in VDC.
Figure 24: Distribution of question type in VDCscore.
19 Calculation of Elo Ranking
Figure 25: GUI screen for Elo ranking.
In this section, we present the methodology for evaluating and ranking AuroraCap and various models using the Elo rating system as shown in Figure 25. The parameters used in the simulation are summarized in Table 20.
  • 1.
    For all models participating in the ELO ranking, we collecte the output captions for each video.
  • 2.
    As shown in Figure 25, we develop a frontend tool to randomly select captions generated by two different models for the same video. Without revealing the model IDs, human evaluators then chose the caption they found better.
  • 3.
    We record comparison results and calculated the ELO values based on the parameters in Table 20.
  • 4.
    In Figure 5, we calculate the Pearson correlation between different metrics and human ELO values, ultimately showing that our proposed VDC metric is the closest to human evaluations.
Table 20: Elo parameter setting.
@@ 19 Calculation of Elo Ranking
20 Case Study
@@
Refer to caption Refer to caption Refer to caption Refer to caption
Figure 26: Video example of MSR-VTT [140] benchmark.
Refer to caption Refer to caption Refer to caption Refer to caption
Figure 27: Video example of VATEX [133] benchmark.
Refer to caption Refer to caption Refer to caption Refer to caption
Figure 28: Video example.
Refer to caption Refer to caption Refer to caption Refer to caption
Figure 29: Video example.
Refer to caption Refer to caption Refer to caption Refer to caption
Figure 30: Video example.
Refer to caption Refer to caption Refer to caption Refer to caption
Figure 31: Video example.
Refer to caption Refer to caption Refer to caption Refer to caption
Figure 32: Video example.
Refer to caption Refer to caption Refer to caption Refer to caption
Figure 33: Video example.
@@
21 Predicted Answer Extraction Prompt Template
Given the question-answer pairs based on the ground truth caption, we utilize GPT-4o to extract predicted answers based on the generated caption by our designed prompt template. The complete prompt is shown as followings:
  • Type
    Prompt
  • SYSTEM
    You are an intelligent chatbot designed for providing accurate answers to questions related to the content based on a detailed description of a video or image.
    Here’s how you can accomplish the task:”
    ##INSTRUCTIONS:
    - Read the detailed description carefully.
    - Answer the question only based on the detailed description.
    - The answer should be a short sentence or phrase.
  • User
    Please provide accurate answers to questions related to the content based on a detailed description of a video or image:
    detailed description: The setting is an urban street with a distinctive red and white painted bike lane, which is part of a larger roadway. The lane is marked with white lines and the word ”WALK” is painted in large, bold letters, indicating a pedestrian crossing area. The pavement is a mix of red brick and grey tiles, with a curb separating the bike lane from the adjacent road. The weather appears to be clear and sunny, casting sharp shadows on the ground, suggesting it is daytime. There are no other vehicles or pedestrians in sight, and the road is empty, which could imply a quiet time of day or a less busy area. The presence of a yellow circular object on the handlebar of the bicycle is notable, possibly a bell or a light.
    question: What objects are the cyclist’s hands interacting with?
    DO NOT PROVIDE ANY OTHER OUTPUT TEXT OR EXPLANATION. Only provide short but accurate answer.
  • GPT-4o
    The cyclist’s hands are interacting with the bicycle handlebars.
22 Correctness Evaluation Prompt Template
Following [87], we evaluate the correctness and score of the predicted answers with the assistant of GPT-4o. Given the question, correct answer, and predicted answer from the generated caption, GPT-4o should return the True or False judgement and relative score (00 to $5$). The complete prompt is shown as followings:
  • Type
    Prompt
  • SYSTEM
    You are an intelligent chatbot designed for evaluating the correctness of generative outputs for question-answer pairs.
    Your task is to compare the predicted answer with the correct answer and determine if they match meaningfully. Here’s how you can accomplish the task:
    ##INSTRUCTIONS:
    - Focus on the meaningful match between the predicted answer and the correct answer.
    - Consider synonyms or paraphrases as valid matches.
    - Evaluate the correctness of the prediction compared to the answer.
  • User
    Please evaluate the following video-based question-answer pair:
    Question: What type of surface can be seen covering the floor?
    Correct Answer: carpet
    Predicted Answer: Carpeted surface.
    Provide your evaluation only as a yes/no and score where the score is an integer value between 0 and 5, with 5 indicating the highest meaningful match.
    Please generate the response in the form of a Python dictionary string with keys ’pred’ and ’score’, where value of ’pred’ is a string of ’yes’ or ’no’ and value of ’score’ is in INTEGER, not STRING.
    DO NOT PROVIDE ANY OTHER OUTPUT TEXT OR EXPLANATION. Only provide the Python dictionary string.
    For example, your response should look like this: {’pred’: ’yes’, ’score’: 4.8}.
  • GPT-4o
    {’pred’: ’yes’, ’score’: 5}